Shown below are the production possibilities curves for two nations: the United States and Chile. Suppose these two nations do not currently engage in international trade or specialization, and suppose that points A and a show the combinations of wheat and copper they now produce and consume United States Chile 10 10 10 10 Copper (thousands of pounds) Copper (thousands of pounds) 1. What is the United States opportunity cost for wheat and copper - show your math work? 2. What is Chiles opportunity cost for wheat and copper -show your math work? 3. Could these two countries gain through specialization? If so, in what products should they each specialize in? 4. Following specialization, suppose the United States exports 3 million of tons of wheat to Chile in exchange for 1 thousand pounds of cooper. What are the gains from trade for each country? (Show your work on how you calculated this number) 5. What are the limits on the terms of trade between the two nations?
Answer #1

Solution:

1. Opportunity cost of United States for wheat = 1/4 copper (thousands of pounds). This is because USA can produce either 8 million tons of wheat or 2 thousands of pounds of copper.

Opportunity cost of United States for copper = 4 wheat (millions of tons).

2. Opportunity cost of Chile for wheat = 1/2 copper (thousands of pounds). This is because Chile can produce either 4 million tons of wheat or 2 thousands of pounds of copper.

Opportunity cost of Chile for copper = 2 wheat (millions of tons).

3. The countries can gain from trade through specialisation. Since USA has absolute advantage in wheat and is at par in respect of copper with Chile, it has comparative advantage in wheat and Chile has comparative advantage in copper.

4. Chile could produce 2 units of wheat in place of 1 unit of copper. If it gets 3 units of wheat from USA, Chile gains 1 unit of wheat from this trade. USA could produce 3/4 unit of copper in place of 3 units of wheat. So gain to USA from trade is (1-3/4)= 0.25 units of copper.

5. Limit on trade is as follows: USA can not exchange more than 4 units of wheat for 1 unit of copper and chile can not exchange less than 2 units of wheat for 1 unit of copper. Hence, 4W>1C>2W is the limit.